If z = 3 - 4i and w = 8 + 3i,write each expression in the standard form a+bi.

A. w - w bar

B. (z - w) bar

Found 2 solutions by ikleyn, math_tutor2020:
Answer by ikleyn(53765)

.

Simple exercise,  which requires the introductory knowledge on complex numbers,  ONLY.

The level of complexity is the same as to add / (to subtract) two integer numbers.

-----------------

On complex numbers,  see introductory lessons
    - Complex numbers and arithmetical operations on them
    - Complex plane
    - Addition and subtraction of complex numbers in complex plane
    - Multiplication and division of complex numbers in complex plane
in this site.



Answer by math_tutor2020(3835)


I'll do the first question to get you started.

w = 8+3i
wBar = complex conjugate of w
wBar = horizontal line over "8+3i" = 8-3i

w - wBar = (8+3i) - (8-3i)
w - wBar = 8+3i - 8+3i
w - wBar = 6i

Notice that in general (a+bi)-(a-bi) = 2bi